OpenStudy (anonymous):

find the x quardinate the y quardinate and equation of the line of semetry -2x^2+2x+2

OpenStudy (anonymous):

-2x^2+2x+2 it dont put it on right

OpenStudy (saifoo.khan):

O ok.

OpenStudy (anonymous):

i have the worse smelling dog he wont stop farting lol... pitbull= bad gas

OpenStudy (saifoo.khan):

Haha.

OpenStudy (saifoo.khan):

Max = y =5/2 at x = 1/2

OpenStudy (saifoo.khan):

y intercpt whn x = 0 -2x^2+2x+2 -2(0)^2+2(0)+2 equals 2

OpenStudy (saifoo.khan):

x co: 1.61 -0.61

OpenStudy (anonymous):

here do this one only like 10 to go..

OpenStudy (anonymous):

x^2-5x+5=0 jeez I dont feel good :(

OpenStudy (saifoo.khan):

please hit good answers!

OpenStudy (saifoo.khan):

3.618 1.382